# Hyperfocal subalgebras of hyperfocal abelian Frobenius blocks

### v1

@@ 6.2 Cyclic group case
Recall from the paragraph 5.16 that we have an indecomposable $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule $M_{\mathrm{st}}$ that induces a stable equivalence of Morita type between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$. Moreover, since the hyperfocal subgroup $D$ is cyclic, $N_{G}(P_{\gamma})$ controls fusion of the block $b$ (see the paragraph 6.16). By Corollary 5.21, the two modules $M_{\mathrm{st}}$ and ${{}_{(u,u)}}M_{\mathrm{st}}$ are isomorphic to each other as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules for any $u\in P$. Fix an isomoprhism $\varphi_{\mathrm{st}}$ between these two bimodules.
Let $\mathbb{P}$ be a projective cover of the bimodule $M_{\mathrm{st}}$ as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules and $\pi$ a corresponding surjective homomorphism from $\mathbb{P}$ to $M_{\mathrm{st}}$. By [10, Theorem 1.1], the hyperfocal subalgebra $\mathbb{A}$ is symmetric. Then by a slight generalization of the proof of [17, Theorem 11.12.1], there is a direct summand $\mathbb{Q}$ of $\mathbb{P}$, uniquely determined by $M_{\mathrm{st}}$ up to isomorphism, such that the following $2$-term complex $\mathbb{M}_{\cdot}$ of $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules
@@
is a Rickard complex of $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules (see the proof of [9, Theorem 5.24]). For simplicity, we still denote $\mathrm{Res}^{\mathbb{P}}_{\mathbb{Q}}(\pi)$ by $\pi$. Obviously, $\pi:{{}_{(u,u)}}\mathbb{P}\rightarrow{{}_{(u,u)}}M_{\mathrm{st}}$ is still a surjective homomorphism of $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules and then $_{(u,u)}$$\mathbb{P}$ is a projective cover of ${{}_{(u,u)}}M_{\mathrm{st}}$. Due to $M_{\mathrm{st}}$ isomorphic to ${{}_{(u,u)}}M_{\mathrm{st}}$ and the uniqueness of projective covers, there is an isomorphism $\psi_{\mathrm{st}}$ of $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules between $\mathbb{P}$ and $_{(u,u)}$$\mathbb{P}$ such that the following diagram commutes
@@
By the construction of $\mathbb{P}$ (see [16, Theorem 11.9.1]), it is clear that every indecomposable direct summand of $\mathbb{P}$ has multiplicity one and $\mathbb{P}$ is isomorphic to $(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})\mathfrak{f}$ as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules for some idempotent $\mathfrak{f}$ of $\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ}$. We can write $_{(u,u)}$$\mathbb{P}$ as $_{(u,u)}$$\mathbb{Q}\oplus\mathbb{Q}^{\prime}$ for some projective $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule $\mathbb{Q}^{\prime}$. By Lemma 6.18, $\mathbb{Q}$ is isomorphic to $_{(u,u)}$$\mathbb{Q}$ as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules. Then we have $\psi_{\mathrm{st}}(\mathbb{Q})$ isomorphic to $_{(u,u)}$$\mathbb{Q}$ as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules. Therefore, by [29, Lemma 44.7], we have ${{}_{(u,u)}}\mathbb{P}=\psi_{\mathrm{st}}(\mathbb{Q})\oplus\mathbb{Q}^{\prime}$ as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules. We can assume that ${{}_{(u,u)}}\mathbb{P}=(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})\mathfrak{f}$. Then the two decompositions ${{}_{(u,u)}}\mathbb{P}={{}_{(u,u)}}\mathbb{Q}\oplus\mathbb{Q}^{\prime}$ and ${{}_{(u,u)}}\mathbb{P}=\psi_{\mathrm{st}}(\mathbb{Q})\oplus\mathbb{Q}^{\prime}$ yields two orthogonal decompositions of the idempotent $\mathfrak{i}$ as follows
@@
with ${{}_{(u,u)}}\mathbb{Q}=(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})\mathfrak{f}_{u}$ and $\psi_{\mathrm{st}}(\mathbb{Q})=(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})\mathfrak{f}_{\mathrm{st}}$ and $(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})\mathfrak{f}_{u}^{\prime}=\mathbb{Q}^{\prime}=(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})\mathfrak{f}_{\mathrm{st}}^{\prime}$. It is easy to check that $\mathfrak{f}_{u}\mathfrak{f}_{\mathrm{st}}^{\prime}=0=\mathfrak{f}_{\mathrm{st}}^{\prime}\mathfrak{f}_{u}$ and $\mathfrak{f}_{\mathrm{st}}\mathfrak{f}_{u}^{\prime}=0=\mathfrak{f}_{u}^{\prime}\mathfrak{f}_{\mathrm{st}}$. Then
$\mathfrak{f}_{u}=\mathfrak{f}\mathfrak{f}_{u}=\mathfrak{f}_{\mathrm{st}}\mathfrak{f}_{u}=\mathfrak{f}_{\mathrm{st}}(\mathfrak{f}_{u}+\mathfrak{f}_{u}^{\prime})=\mathfrak{f}_{\mathrm{st}}.$
@@
6.22.
Next, we will investigate the structures of $\mathrm{Irr}_{\mathcal{K}}(A)$ and $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})$. Set $|D|=p^{n}$ with $n\geq 1$ and $|E|=e$. Clearly, $e$ divides p1p-1. By Theorem 6.20, $\mathrm{k}(\mathbb{A})=|\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})|=|\mathrm{Irr}(L)|=e+\frac{p^{n}-1}{e}$ and $l(\mathbb{A})=|\mathrm{Irr}_{k}(\mathbb{A})|=|\mathrm{IBr}(L)|=e$. Denote by $\mathcal{M}$ a set of representatives of the $E$-conjugacy classes of nontrivial irreducible characters of $D$. Clearly, $|\mathcal{M}|=\frac{p^{n}-1}{e}$. We can identify $\mathrm{Irr}(E)$ with a subset of $\mathrm{Irr}(L)$ through the canonical way. Then $\mathrm{Irr}(L)=\{\mathrm{Ind}_{D}^{L}(\chi)\mid\chi\in\mathcal{M}\}\cup\{\lambda\mid\lambda\in\mathrm{Irr}(E)\}$. So we can set $\{\hat{U}_{\chi},\hat{U}_{\lambda}\mid\chi\in\mathcal{M},\lambda\in\mathrm{Irr}(E)\}$ to be a set of representatives of the isomorphism classes of simple $\hat{\mathbb{A}}-$modules and $\{\bar{W}_{\lambda}\mid\lambda\in\mathrm{Irr}(E)\}$ to be a set of representatives of the isomorphism classes of simple $\bar{\mathbb{A}}$-modules. By [12, Proposition 3.2], the decomposition map $d_{\mathbb{A}}$ is surjective. At the same time, by Corollary 5.15, the Cartan matrix of $\mathbb{A}$ is non-singular. Therefore, by [12, Proposition 2.2], we have $L^{0}(\mathbb{A})^{\bot}=\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$. For the group algebra $\mbox{$\mathcal{O}$}L$, we denote by $R_{\mathcal{K}}(\mbox{$\mathcal{O}$}L)$ the set of all generalized characters of $L$. We adopt the similar notation for $R_{k}(\mbox{$\mathcal{O}$}L)$.
6.23.
@@
Then the restricting of $\Phi_{\mathcal{K}}$ to $L^{0}(\mathbb{A})$ induces a bijective isometry, still denoted by $\Phi_{\mathcal{K}}$ between $L^{0}(\mathbb{A})$ and $L^{0}(\mbox{$\mathcal{O}$}L)$. Furthermore, since $\Phi_{\mathcal{K}}$ is an isometry, we can get an bijective isometry $\Phi_{\mathcal{K}}:\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})\rightarrow\mathrm{Pr}_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)$. Then we can obtain $\mathbb{Z}$-bases of $\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$ and $L^{0}(\mathbb{A})$ as follows.
@@
(iii) $\{\epsilon_{\lambda}[\hat{U}_{\lambda}]+\sum\limits_{\chi\in\mathcal{M}}\epsilon_{\chi}[\hat{U}_{\chi}]\mid\lambda\in\mathrm{Irr}(E)\}$ is a $\mathbb{Z}$-basis of $\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$.
@@
Let $\sum\limits_{\chi\in\mathcal{M},\lambda\in\mathrm{Irr}(E)}(m_{\chi}[\hat{U}_{\chi}]+m_{\lambda}[\hat{U}_{\lambda}])$ be an element of $L^{0}(\mathbb{A})$. Here, $m_{\chi}$ and $m_{\lambda}$ are integers. We have $\sum\limits_{\chi\in\mathcal{M},\lambda\in\mathrm{Irr}(E)}(m_{\chi}\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\chi})+m_{\lambda}\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\lambda}))=0$.
@@
It can be easily deduced from the well-known fact that $\mathrm{dim}_{\mathcal{K}}(\mathrm{Hom}_{\hat{\mathbb{A}}}(\hat{\mathbb{A}}\mathfrak{i},X))=\mathrm{dim}_{\mathcal{K}}(\mathfrak{i}X)$ for any primitive idempotent $\mathfrak{i}$ of $\mathbb{A}$ and any finite dimensional $\hat{\mathbb{A}}-$module $X$. ∎
6.26.
Now we can verify the forward direction of the KLN conjecture when the hyperfocal subgroup $D$ is nontrivial cyclic. We first recall a well-known fact that there is a simple $\hat{A}-$module of dimension coprime to $p$. Indeed, by [17, Proposition 6.11.11], the $p$-part of the dimension of a simple $\hat{A}-$module equals to the height of its corresponding irreducible ordinary character in the block $b$ through the Morita equivalence induced by the bimodule $i\mbox{$\mathcal{O}$}G$.
@@
This can be easily deduced from the argument above and [13, Proposition 2.3 (iv)]. ∎
@@
Since the hyperfocal subgroup $D$ is assumed to be nontrivial, the order $e$ of $E$ can not be $1$. Then by Proposition 6.24 (iii), $\epsilon_{\lambda}[\hat{U}_{\lambda}]-\epsilon_{\mu}[\hat{U}_{\mu}]$ belongs to $\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$ for any two distinct elements $\lambda$ and $\mu$ of $\mathrm{Irr}(E)$. Since $\mathbb{A}$ is projective as left $\mbox{$\mathcal{O}$}D$-module, we have $p$ dividing $\epsilon_{\lambda}\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\lambda})-\epsilon_{\mu}\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\mu})$. In particular, $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\lambda})$ is coprime to $p$ if and only if $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\mu})$ is coprime to $p$. Suppose that $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\lambda})$ is divided by $p$ for some $\lambda\in\mathrm{Irr}(E)$. So is $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\mu})$ for any $\mu\in\mathrm{Irr}(E)$. Now by Proposition 6.24 (ii) and Lemma 6.25, $\epsilon\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\chi})-\sum\limits_{\lambda\in\mathrm{Irr}(E)}\epsilon_{\lambda}\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\lambda})=0$. Hence, $\epsilon\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\chi})$ is divided by $p$ for any $\chi\in\mathcal{M}$. In conclusion, every simple $\hat{\mathbb{A}}-$module has dimension divided by $p$, which contradicts Lemma 6.27. So $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\lambda})$ is coprime to $p$ for any $\lambda\in\mathrm{Irr}(E)$. By Proposition 6.24 (iii), there is at least one element $\chi$ of $\mathcal{M}$ such that $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\chi})$ is coprime to $p$. We denote it by $\chi_{0}$. Now we can assume that $\mathcal{M}$ has at least two elements. For any other $\chi\in\mathcal{M}$ different from $\chi_{0}$, by Proposition 6.24 (ii), $\epsilon([\hat{U}_{\chi_{0}}]-[\hat{U}_{\chi}])$ belongs to $L^{0}(\mathbb{A})$. In particular, by Lemma 6.25, these two simple modules $\hat{U}_{\chi_{0}}$ and $\hat{U}_{\chi}$ has the same dimension. Then we are done. ∎
6.29.
@@
Note that the structure of a $\mathbb{Z}$-basis of $\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$ in Proposition 6.24 plays an important role in the proof of Proposition 6.28. Furthermore, by this structure, we can also calculate decomposition numbers of $\mathbb{A}$ and get that these numbers are either $1$ or 0 in the following. This result can be regarded as a ‘hyperfocal decomposition number’ version of the classical fact that decomposition numbers of blocks with cyclic defect groups are either $1$ or 0. By [12, §2], this is equivalent to showing that $\hat{\mathbb{A}}\mathfrak{i}]=\sum\limits_{\theta\in\mathrm{Irr}(L)}a_{\theta}[\hat{U}_{\theta}]$ with $a_{\theta}$ equal to $1$ or 0 for any primitive idempotent $\mathfrak{i}$ of $\mathbb{A}$. One key technique in the context of blocks with cyclic defect groups is making use of $\mathcal{O}$-pure submodules. An $\mathbb{A}$-submodule $U^{\prime}$ of an $\mathbb{A}$-module $U$ is $\mathcal{O}$-pure in $U$ if it is a direct summand of $U$ as an $\mathcal{O}$-module (see [16, Section 4.2]). Our proof is similar to the one of [17, Theorem 11.10.5]. We need the following two lemmas which are analogues of [17, Lemma 11.10.3] and [17, Proposition 11.10.4].
@@
The key ingredient in the proof of [17, Lemma 11.10.3] is the existence of a stable equivalence induced by induction and restriction. Recall from Section $5$ that there is an $\mathcal{O}$-algebra $\mathbb{A}^{\prime}$ Morita equivalent to $\mathbb{A}$ such that it has a subalgebra isomorphic to $\mbox{$\mathcal{O}$}L$ and a stable equivalence of Morita type between $\mathrm{A}^{\prime}$ and this subalgebra induced by induction and restriction exists. Then the proof of [17, Lemma 11.10.3] can be applied to our case. ∎

### v2

@@ 6.2 Cyclic group case
Recall from the paragraph 5.16 that we have an indecomposable $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule $M_{\mathrm{st}}$ that induces a stable equivalence of Morita type between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$. By Proposition 5.19, the two modules $M_{\mathrm{st}}$ and ${{}_{(u,u)}}M_{\mathrm{st}}$ are isomorphic to each other as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules for any $u\in P$. Fix an isomoprhism $\varphi_{\mathrm{st}}$ between these two bimodules.
Let $\mathbb{P}$ be a projective cover of the bimodule $M_{\mathrm{st}}$ as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules and $\pi$ a corresponding surjective homomorphism from $\mathbb{P}$ to $M_{\mathrm{st}}$. By [10, Theorem 1.1], the hyperfocal subalgebra $\mathbb{A}$ is symmetric. Then by a slight generalization of the proof of [16, Theorem 11.12.1], there is a direct summand $\mathbb{Q}$ of $\mathbb{P}$, uniquely determined by $M_{\mathrm{st}}$ up to isomorphism, such that the following $2$-term complex $\mathbb{M}_{\cdot}$ of $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules
@@
is a Rickard complex of $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules (see the proof of [9, Theorem 5.24]). For simplicity, we still denote $\mathrm{Res}^{\mathbb{P}}_{\mathbb{Q}}(\pi)$ by $\pi$. Obviously, $\pi:{{}_{(u,u)}}\mathbb{P}\rightarrow{{}_{(u,u)}}M_{\mathrm{st}}$ is still a surjective homomorphism of $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules and then $_{(u,u)}$$\mathbb{P}$ is a projective cover of ${{}_{(u,u)}}M_{\mathrm{st}}$. Due to the facts that $M_{\mathrm{st}}$ is isomorphic to ${{}_{(u,u)}}M_{\mathrm{st}}$ and projective covers are unique up to isomorphism, there is an isomorphism $\psi_{\mathrm{st}}$ of $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules between $\mathbb{P}$ and $_{(u,u)}$$\mathbb{P}$ such that the following diagram commutes
@@
By the construction of $\mathbb{P}$ (see [15, Theorem 11.9.1]), it is clear that every indecomposable direct summand of $\mathbb{P}$ has multiplicity one and $\mathbb{P}$ is isomorphic to $(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})\mathfrak{f}$ as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules for some idempotent $\mathfrak{f}$ of $\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ}$. We can write $_{(u,u)}$$\mathbb{P}$ as $_{(u,u)}$$\mathbb{Q}\oplus\mathbb{Q}^{\prime}$ for some projective $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule $\mathbb{Q}^{\prime}$. By Lemma 6.20, $\mathbb{Q}$ is isomorphic to $_{(u,u)}$$\mathbb{Q}$ as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules. Then we have $\psi_{\mathrm{st}}(\mathbb{Q})$ isomorphic to $_{(u,u)}$$\mathbb{Q}$ as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules. Therefore, by [29, Lemma (44.7)], we have ${{}_{(u,u)}}\mathbb{P}=\psi_{\mathrm{st}}(\mathbb{Q})\oplus\mathbb{Q}^{\prime}$ as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules. We can assume that ${{}_{(u,u)}}\mathbb{P}=(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})\mathfrak{f}$. Then the two decompositions ${{}_{(u,u)}}\mathbb{P}={{}_{(u,u)}}\mathbb{Q}\oplus\mathbb{Q}^{\prime}$ and ${{}_{(u,u)}}\mathbb{P}=\psi_{\mathrm{st}}(\mathbb{Q})\oplus\mathbb{Q}^{\prime}$ yields two orthogonal decompositions of the idempotent $\mathfrak{f}$ as follows
@@
with ${{}_{(u,u)}}\mathbb{Q}=(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})\mathfrak{f}_{u}$ and $\psi_{\mathrm{st}}(\mathbb{Q})=(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})\mathfrak{f}_{\mathrm{st}}$ and $(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})\mathfrak{f}_{u}^{\prime}=\mathbb{Q}^{\prime}=(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})\mathfrak{f}_{\mathrm{st}}^{\prime}$. By the equation $(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})\mathfrak{f}_{u}^{\prime}=(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})\mathfrak{f}_{\mathrm{st}}^{\prime}$, we can get that $\mathfrak{f}^{\prime}_{u}\mathfrak{f}^{\prime}_{{\rm st}}=\mathfrak{f}^{\prime}_{u}$ and $\mathfrak{f}^{\prime}_{{\rm st}}\mathfrak{f}^{\prime}_{u}=\mathfrak{f}^{\prime}_{{\rm st}}$. Then it is easy to check that $\mathfrak{f}_{u}\mathfrak{f}_{\mathrm{st}}^{\prime}=0=\mathfrak{f}_{\mathrm{st}}^{\prime}\mathfrak{f}_{u}$ and $\mathfrak{f}_{\mathrm{st}}\mathfrak{f}_{u}^{\prime}=0=\mathfrak{f}_{u}^{\prime}\mathfrak{f}_{\mathrm{st}}$. Hence,
$\mathfrak{f}_{u}=\mathfrak{f}\mathfrak{f}_{u}=\mathfrak{f}_{\mathrm{st}}\mathfrak{f}_{u}=\mathfrak{f}_{\mathrm{st}}(\mathfrak{f}_{u}+\mathfrak{f}_{u}^{\prime})=\mathfrak{f}_{{\rm st}}\mathfrak{f}=\mathfrak{f}_{\mathrm{st}}.$
@@
6.24.
Next, we will investigate the structures of $\mathrm{Irr}_{\mathcal{K}}(A)$ and $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})$. Set $|D|=p^{n}$ with $n\geq 1$ and $|E|=e$. Clearly, $e$ divides p1p-1. By Theorem 6.22, $\mathrm{k}(\mathbb{A})=|\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})|=|\mathrm{Irr}(L)|=e+\frac{p^{n}-1}{e}$ and $l(\mathbb{A})=|\mathrm{Irr}_{k}(\mathbb{A})|=|\mathrm{IBr}(L)|=e$. Denote by $\mathcal{M}$ a set of representatives of the $E$-conjugacy classes of nontrivial irreducible characters of $D$. Clearly, $|\mathcal{M}|=\frac{p^{n}-1}{e}$. We can identify $\mathrm{Irr}(E)$ with a subset of $\mathrm{Irr}(L)$ through the canonical way. Then $\mathrm{Irr}(L)=\{\mathrm{Ind}_{D}^{L}(\chi)\mid\chi\in\mathcal{M}\}\cup\{\lambda\mid\lambda\in\mathrm{Irr}(E)\}$. So by the Rickard equivalence obtained in Theorem 6.22, we can set $\{\hat{U}_{\chi},\hat{U}_{\lambda}\mid\chi\in\mathcal{M},\lambda\in\mathrm{Irr}(E)\}$ to be a set of representatives of the isomorphism classes of simple $\hat{\mathbb{A}}-$modules and $\{\bar{W}_{\lambda}\mid\lambda\in\mathrm{Irr}(E)\}$ to be a set of representatives of the isomorphism classes of simple $\bar{\mathbb{A}}$-modules. By [11, Proposition 3.2], the decomposition map $d_{\mathbb{A}}$ is surjective. At the same time, by Corollary 5.15, the Cartan matrix of $\mathbb{A}$ is non-singular. Therefore, by [11, Proposition 2.2], we have $L^{0}(\mathbb{A})^{\bot}=\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$. For the group algebra $\mbox{$\mathcal{O}$}L$, we denote by $R_{\mathcal{K}}(\mbox{$\mathcal{O}$}L)$ the set of all generalized characters of $L$. We adopt the similar notation for $R_{k}(\mbox{$\mathcal{O}$}L)$.
6.25.
@@
Then the restricting of $\Phi_{\mathcal{K}}$ to $L^{0}(\mathbb{A})$ induces a bijective isometry, still denoted by $\Phi_{\mathcal{K}}$, between $L^{0}(\mathbb{A})$ and $L^{0}(\mbox{$\mathcal{O}$}L)$. Furthermore, since $\Phi_{\mathcal{K}}$ is an isometry, we can get an bijective isometry $\Phi_{\mathcal{K}}:\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})\rightarrow\mathrm{Pr}_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)$. Then we can obtain $\mathbb{Z}$-bases of $\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$ and $L^{0}(\mathbb{A})$ as follows.
@@
(iii) $\{\epsilon_{\lambda}[\hat{U}_{\lambda}]+\epsilon\sum\limits_{\chi\in\mathcal{M}}[\hat{U}_{\chi}]\mid\lambda\in\mathrm{Irr}(E)\}$ is a $\mathbb{Z}$-basis of $\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$.
@@
Let $\sum\limits_{\chi\in\mathcal{M},\lambda\in\mathrm{Irr}(E)}(m_{\chi}[\hat{U}_{\chi}]+m_{\lambda}[\hat{U}_{\lambda}])$ be an element in $L^{0}(\mathbb{A})$. Here, $m_{\chi}$ and $m_{\lambda}$ are integers. We have $\sum\limits_{\chi\in\mathcal{M},\lambda\in\mathrm{Irr}(E)}(m_{\chi}\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\chi})+m_{\lambda}\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\lambda}))=0$.
@@
It is clear that taking the dimension of any simple $\hat{\mathbb{A}}-$module can induce a $\mathbb{Z}$-linear map ${\rm dim}_{\mathcal{K}}$ from $R_{\mathcal{K}}(\mathbb{A})$ to $\mathbb{Z}$. Similarly, taking the dimension of any simple $\bar{\mathbb{A}}$-module can also induce a $\mathbb{Z}$-linear map ${\rm dim}_{k}$ from $R_{k}(\mathbb{A})$ to $\mathbb{Z}$. Let $X$ be a finite dimensional $\hat{\mathbb{A}}-$module. Set $d_{\mathbb{A}}([X])=\sum\limits_{\lambda\in{\rm Irr}(E)}a_{\lambda}[\bar{W}_{\lambda}]$ for some nonnegative integer $a_{\lambda}$. By the definition of the decomposition map $d_{\mathbb{A}}$, we can get that ${\rm dim}_{\mathcal{K}}(X)=\sum\limits_{\lambda\in{\rm Irr}(E)}a_{\lambda}{\rm dim}_{k}(\bar{W}_{\lambda})$. This implies that ${\rm dim}_{k}\circ d_{\mathbb{A}}={\rm dim}_{\mathcal{K}}$. Since $L^{0}(\mathbb{A})$ is the kernel of the decomposition map $d_{\mathbb{A}}$, we complete the proof of this lemma. ∎
6.28.
Now we can verify the forward direction of the KLN conjecture when the hyperfocal subgroup $D$ is nontrivial cyclic. We first recall a well-known fact that there is a simple $\hat{A}-$module of dimension coprime to $p$. Indeed, by [16, Proposition 6.11.11], the $p$-part of the dimension of a simple $\hat{A}-$module equals to the height of its corresponding irreducible ordinary character in the block $b$ through the Morita equivalence induced by the bimodule $i\mbox{$\mathcal{O}$}G$.
@@
This can be easily deduced from the argument above and [12, Proposition 2.3 (iv)]. ∎
@@
Since the hyperfocal subgroup $D$ is assumed to be nontrivial, the order $e$ of $E$ can not be $1$. Then by Proposition 6.26 (iii), $\epsilon_{\lambda}[\hat{U}_{\lambda}]-\epsilon_{\mu}[\hat{U}_{\mu}]$ belongs to $\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$ for any two distinct elements $\lambda$ and $\mu$ of $\mathrm{Irr}(E)$. Since $\mathbb{A}$ is projective as left $\mbox{$\mathcal{O}$}D$-module, every finitely generated projective $\mathbb{A}$-module has $\mathcal{O}$-rank divided by $|D|$. Hence, we have $p$ dividing $\epsilon_{\lambda}\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\lambda})-\epsilon_{\mu}\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\mu})$. In particular, $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\lambda})$ is coprime to $p$ if and only if $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\mu})$ is coprime to $p$. Suppose that $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\lambda})$ is divided by $p$ for some $\lambda\in\mathrm{Irr}(E)$. So is $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\mu})$ for any $\mu\in\mathrm{Irr}(E)$. Now by Proposition 6.26 (ii) and Lemma 6.27, $\epsilon\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\chi})-\sum\limits_{\lambda\in\mathrm{Irr}(E)}\epsilon_{\lambda}\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\lambda})=0$. Hence, $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\chi})$ is divided by $p$ for any $\chi\in\mathcal{M}$. In conclusion, every simple $\hat{\mathbb{A}}-$module has dimension divided by $p$, which contradicts Lemma 6.29. So $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\lambda})$ is coprime to $p$ for any $\lambda\in\mathrm{Irr}(E)$. By Proposition 6.26 (iii), there is at least one element $\chi$ in $\mathcal{M}$ such that $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\chi})$ is coprime to $p$. We denote it by $\chi_{0}$. Now we can assume that $\mathcal{M}$ has at least two elements. For any other $\chi\in\mathcal{M}$ different from $\chi_{0}$, by Proposition 6.26 (ii), $\epsilon([\hat{U}_{\chi_{0}}]-[\hat{U}_{\chi}])$ belongs to $L^{0}(\mathbb{A})$. In particular, by Lemma 6.27, these two simple modules $\hat{U}_{\chi_{0}}$ and $\hat{U}_{\chi}$ has the same dimension. Then we are done. ∎
6.31.
@@
Note that the structure of a $\mathbb{Z}$-basis of $\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$ in Proposition 6.26 plays an important role in the proof of Proposition 6.30. Furthermore, by this structure, we can also calculate decomposition numbers of $\mathbb{A}$ and get that these numbers are either $1$ or 0 in the following. This result can be regarded as a ‘hyperfocal decomposition number’ version of the classical fact that decomposition numbers of blocks with nontrivial cyclic defect groups are either $1$ or 0. By [11, §2], this is equivalent to showing that $\mathcal{K}\mathop{\otimes}\limits_{\mathcal{O}}U]=\sum\limits_{\theta\in\mathrm{Irr}(L)}a_{\theta}[\hat{U}_{\theta}]$ with $a_{\theta}$ equal to $1$ or 0 for any projective indecomposable $\mathbb{A}$-module $U$. One key technique in the context of blocks with nontrivial cyclic defect groups is making use of $\mathcal{O}$-pure submodules. An $\mathbb{A}$-submodule $U^{\prime}$ of an $\mathbb{A}$-module $U$ is $\mathcal{O}$-pure in $U$ if it is a direct summand of $U$ as $\mathcal{O}$-modules (see [15, Section 4.2]). Our proof here is similar to the one of [16, Theorem 11.10.5]. We need the following two lemmas which are analogues of [16, Lemma 11.10.3] and [16, Proposition 11.10.4].
@@
[…]